OpenStudy (anonymous):

find the sum of 2 radical 6 and radical 24

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow \sqrt{24} = \sqrt{2 * 2 * 2 * 3} }\) \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{24} = 2\sqrt{6} }\) Think of x as root 6. What is 2x + 2x? 4x, so: \(\Large \color{MidnightBlue}{\Rightarrow 4\sqrt{6} }\)